OpenStudy (anonymous):

Three equal resistors each of 12ohm can be connected in four different ways. What is equivalent resistance of each combination?

OpenStudy (anonymous):

hi can you figure ways ?

OpenStudy (nikvist):

\[R_1=R_2=R_3=R=12\Omega\]\[R_{12S}=2R=24\Omega\quad,\quad R_{12P}=R/2=6\Omega\]\[1.\quad R_{E1}=R_{12S}+R_3=3R=36\Omega\]\[2.\quad R_{E2}=R_{12S}\parallel R_3=\frac{R_{12S}R_3}{R_{12S}+R_3}=\frac{2R}{3}=8\Omega\]\[3.\quad R_{E3}=R_{12P}+R_3=\frac{3R}{2}=18\Omega\]\[4.\quad R_{E4}=R_{12P}\parallel R_3=\frac{R_{12P}R_3}{R_{12P}+R_3}=\frac{R}{3}=4\Omega\]